OpenStudy (anonymous):

how do you convert miles per hour to feet per second? Specifically I need to find 233mi/hr in ft/sec

OpenStudy (anonymous):

use dimensional analysis

OpenStudy (anonymous):

What do you mean?

OpenStudy (dumbcow):

\[\frac{X mi}{hr} * \frac{1 hr}{3600 \sec}*\frac{5280 ft}{1 mi}\]

OpenStudy (anonymous):

That helps:)

OpenStudy (anonymous):

Ok but where would the numbers fit? I'm confused sorry

OpenStudy (dumbcow):

that reduces to 22/15 so the conversion is \[1 m/hr = \frac{22}{15} ft/\sec\]

OpenStudy (anonymous):

So I have to put in 233 into the equation to solve it right?

OpenStudy (dumbcow):

right... 233 mi/hr = 233*(22/15) ft/sec

OpenStudy (anonymous):

Um i got 5126/3495...I have to simplify it so

OpenStudy (dumbcow):

what happened to your denominator ? it should be 5126/15

OpenStudy (dumbcow):

i don't know, round to 2 decimal places i guess

OpenStudy (anonymous):

huh?ohhhhhhhhh!!!! Wow!!

OpenStudy (anonymous):

that is already the answer

OpenStudy (anonymous):

Just use dumbcows formula and plug in your number in the X the rest should cancle out and be left with a coeffecient and ft/s

OpenStudy (anonymous):

so 233mi/hr = 5126/15ft/sec?

OpenStudy (anonymous):

yes , I think there is no difference if that is your answer or if you do not want fraction your answer will be 341.73 ft/sec.

OpenStudy (anonymous):

oh ok that makes sense:) Thanks to everyone! Sorry I was a bit drowsy:/